OpenStudy (anonymous):

Please place the following in y=mx+b format. No decimals please also place all fractions as improper and in lowest terms.Also please explan how you got the answer.

OpenStudy (anonymous):

the problem is: y-7=1/4(x-3)

OpenStudy (anonymous):

first multiply out using the distributive law to get \[y-7=\frac{1}{4}x-\frac{3}{4}\] then add 7 to both sides to get \[y=\frac{1}{4}x+\frac{25}{4}\]

OpenStudy (anonymous):

Thank you lots